OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1. Show your work for full credit.

OpenStudy (dumbcow):

\[\frac{y}{x} = \tan \theta\] \[\tan (-\frac{3\pi}{4}) = 1\] y = -1 point: (-1,-1)

OpenStudy (anonymous):

how did you get y=-1?

OpenStudy (dumbcow):

x is given as -1 and tangent = 1 y/(-1) = 1 --> y = -1

OpenStudy (anonymous):

oh! thank you a lot :) have a good day